• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 00:07:26 on February 06, 2009

    Entry number 259133


    keyword=Start/End Run Script Status for d2n

    The start and end run scripts have been updated for the d2n experiment to
    run with two independent DAQs. For the BigBite side running on adaql2, I
    made only a few minor changes such as dropping the RICH and A1 EPICS
    variables. For the left HRS DAQ running on adaql1, I basically copied all
    the relevant files that we had been using for Transversity on adaql2 to
    adaql1. The way these files are set up, all the paths are reference to
    the $HOME directory on which the script is run. Hence, coping the
    scripts over should not cause the left HRS DAQ to step on the BigBite DAQ.

    I will go through the left HRS scripts closely tomorrow to make sure I
    haven't missed anything.

    What still needs to be done is to get the end-of-run scaler information
    correct in the two DAQs. Currently neither DAQ has the right
    information in the end-of-run file.

    The epics loggers on the BigBite side are completed. I still need to do
    the same for the left HRS.

    A copy of this log entry has been emailed to: brads,rom,moffit,camsonne